OpenStudy (anonymous):

Simplest form 4^-3.5

OpenStudy (anonymous):

4^-3.5 When a negative exponent ya flip so it would be \[\frac{ 1 }{ 4^{3.5} }\] 4^3.5 can be rewritten as\[4^{\frac{ 7 }{ 2 }} meaning \rightarrow \sqrt[2]{4}^{7}\]

OpenStudy (anonymous):

square of 4 is 2 and 2^7 is

OpenStudy (anonymous):



OpenStudy (anonymous):

yea i know your finally answer would be 1/128

OpenStudy (anonymous):

When you have a negative exponent you flip it

OpenStudy (anonymous):

wow ok thanks . i have another

OpenStudy (anonymous):

ok

OpenStudy (anonymous):